OpenStudy (anonymous):

find the length of the curve y=(9-x^2/3)^3/2 from x=1 to x=27

OpenStudy (anonymous):

Integrate buddy.

OpenStudy (anonymous):

\[L=\int\limits \sqrt{1+(\frac{dy}{dx})^2}dx\], right?

OpenStudy (anonymous):

Pythagorean distance formula..

OpenStudy (shayaan_mustafa):

Hi roj3og :) How are you dear? Its arc length. Right?

OpenStudy (shayaan_mustafa):

@CliffSedge integrate from a to b. lower limit is a and upper is b

OpenStudy (anonymous):

yeah that is what the book says to do but I can't seem to find the right dy/dx arc length correct

OpenStudy (anonymous):

@Shayaan_Mustafa , yes, I know what limits are..

OpenStudy (shayaan_mustafa):

Why? It is so simple . Just differentiate w.r.t. to x. Is it difficult my friend?

OpenStudy (anonymous):

How far along did you get? Were you able to find dy/dx? Do you know what function you need to integrate?

OpenStudy (anonymous):

is this correct for the dy/dx \[-((9-x ^{2/3})/x ^{1/3} ) ^{1/2}\]

OpenStudy (anonymous):

not sure where the x^1/3 came from..

OpenStudy (anonymous):

or the x^2/3..

OpenStudy (anonymous):

Use chain rule. \[\frac{d}{dx} u(x)^{3/2} = \frac{3}{2}u(x)^{1/2}u'(x) : u=9-\frac{x^2}{3}\]

OpenStudy (anonymous):

n.b. u(x)^3/2 means the function, u-of-x raised to 3/2, not u times x to the 3/2.

OpenStudy (anonymous):

Am I reading your function incorrectly? Is it\[y=(9-\frac{x^2}{3})^{3/2}\] or \[y=(9-x^{2/3})^{3/2}?\]

OpenStudy (anonymous):

If it's the latter, then your derivative was almost correct. \[y=(9-x^{2/3})^{3/2} \rightarrow \frac{dy}{dx}=-\frac{(9-x^{2/3})^{1/2}}{x^{1/3}}\]

OpenStudy (anonymous):

it is the second one.

OpenStudy (anonymous):

Were you able to do the integration after finding that derivative?

OpenStudy (anonymous):

\[ 1+y'(x)^2=\frac{9-x^{2/3}}{x^{2/3}}+1=\frac{9}{x^{2/3}} \]

OpenStudy (anonymous):

\[ \sqrt{1+y'(x)^2}=\sqrt{\frac{9}{x^{2/3}}} =\frac{3}{\sqrt[3]{x}} \]

OpenStudy (anonymous):

eliassaab, thank you but that is what i finally came up with as well and the homework doesn't give that as an answer. the options they give me are. 28/3, 14/3, 21, or 14

OpenStudy (anonymous):

\[ \int_1^{27} \frac{3}{\sqrt[3]{x}} \, dx=36 \]

OpenStudy (anonymous):

Either you typed the problem wrong or the choices are not correct.

OpenStudy (anonymous):

I was looking at the answers for another problem, HAHA. eliassaab you saved me on that one. I've been working on this for too long, everything is starting to blur together.

OpenStudy (anonymous):

thanks everyone!